# F-Bench: Rethinking Human Preference Evaluation Metrics for Benchmarking Face Generation, Customization, and Restoration

### v1

@@ Abstract
Artificial intelligence generative models exhibit remarkable capabilities in content creation, particularly in face image generation, customization, and restoration. However, current AI-generated faces (AIGFs) often fall short of human preferences due to unique distortions, unrealistic details, and unexpected identity shifts, underscoring the need for a comprehensive quality evaluation framework for AIGFs. To address this need, we introduce FaceQ, a large-scale, comprehensive database of AI-generated Face images with fine-grained Quality annotations reflecting human preferences. The FaceQ database comprises 12,255 images generated by 29 models across three tasks: (1) face generation, (2) face customization, and (3) face restoration. It includes 32,742 mean opinion scores (MOSs) from 180 annotators, assessed across multiple dimensions: quality, authenticity, identity (ID) fidelity, and text-image correspondence. Using the FaceQ database, we establish F-Bench, a benchmark for comparing and evaluating face generation, customization, and restoration models, highlighting strengths and weaknesses across various prompts and evaluation dimensions. Additionally, we assess the performance of existing image quality assessment (IQA), face quality assessment (FQA), AI-generated content image quality assessment (AIGCIQA), and preference evaluation metrics, manifesting that these standard metrics are relatively ineffective in evaluating authenticity, ID fidelity, and text-image correspondence. The FaceQ database will be publicly available upon publication.
@@ 1 Introduction
Generative models, including VAE [34], GAN [17], and diffusion models [25, 77, 66], have achieved remarkable advancements in recent years. Numerous variants of these foundational models leverage generative priors for various face-related downstream tasks, such as face generation [68, 71, 63], face customization [44, 103], and face restoration [49, 84], producing creative, high-quality, and realistic face images.
However, despite these advancements, current AI-generated faces (AIGFs) still exhibit a range of perceptual flaws. Unlike natural face images, which may suffer from low-level degradations such as compression artifacts, noise, or blur, AIGFs often lack authenticity in human perception. For example, as shown in Fig.1, unique distortions in facial attributes, including AI-generated artifacts and misalignments, contribute to a notable bad face issue. Unrealistic details, such as over-glossy skin or disproportionate facial features, negatively impact user preference. Additionally, unexpected identity variations frequently arise in face customization and restoration tasks. These prevalent problems highlight the necessity of establishing a comprehensive quality evaluator for AIGFs.
Current face quality assessment databases [3, 79, 7, 50] are primarily based on real face images sourced from photographers or online collections, which differ significantly from AI-generated faces in terms of texture, structure, and artifact characteristics. While recent AIGC quality assessment efforts [35, 96, 38, 83, 46] have focused on general image quality, they pay little attention to the unique challenges posed by face images. Without an AIGF-specific database, existing quality assessment (QA) methods struggle to evaluate AI-generated faces accurately, often failing to address issues like authenticity and identity fidelity, which are central to the user’s perception of face images.
To address this gap, we propose FaceQ, a large-scale database for AI-generated faces, each meticulously annotated with fine-grained human preference scores. As illustrated in Fig.1, the FaceQ database contains 12,255 images generated by 29 generative models across three tasks: 1) face generation, 2) face customization, and 3) face restoration. For face generation and customization, diverse source content is collected, covering various categories of prompts and identities. For face restoration, both synthetic and real-world degradations are included through a mixed, random degradation pipeline. In total, 32,742 mean opinion scores (MOSs) are obtained from 180 annotators across multiple evaluation dimensions, including quality, authenticity, identity (ID) fidelity, and text-image correspondence. Each dimension exhibits distinct distributions and characteristics.
Based on the proposed FaceQ database, we establish a subjective human preference benchmark, F-Bench, by comparing and evaluating current models for face generation, customization, and restoration. Model-wise advantages and disadvantages are thoroughly analyzed across various dimensions for all three tasks. In-depth class-wise comparisons provide a finer-grained evaluation of the models’ performance with respect to prompts and demographic factors. Additionally, we compare the performance of existing image quality assessment (IQA), face quality assessment (FQA), AI-generated content image quality assessment (AIGCIQA), and preference evaluation metrics on their correlation with human, demonstrating that these mainstream metrics perform relatively poor in evaluating the authenticity, identity fidelity, and text-to-image correspondence. Overall, our study provides valuable insights into current face generation, customization, and restoration models, offering guidance for the development of more accurate AIGF evaluators.
@@ 2 Related Work
| Database | Domain | Source | Tasks | Images | Scores | Ratings | Dimensions |
| Pick-A-Pic (NeurIPS2023) [35] | General | AI-generated | 1 | 500000 | Pair | 500000 | Overall |
| HPS (ICCV2023) [96] | General | AI-generated | 1 | 98807 | Pair | 98807 | Overall |
| AGIQA-3K (TCSVT2024) [38] | General | AI-generated | 1 | 2982 | MOS | 5964 | Perception, Alignment |
| AIGCIQA (CVPR2024) [83] | General | AI-generated | 1 | 2400 | MOS | 7200 | Quality, Authenticity, Correspondence |
| RichHF-18K (CVPR2024) [46] | General | AI-generated | 1 | 17760 | MOS | 71040 | Plausibility, Aesthetics, Text-image Alignment, Overall |
[…]
| CFIQA-20k (TMM2023) [79] | Face | Real | 1 | 20000 | MOS | 20000 | Overall |
| CGFIQA-40k (CVPR2024) [7] | Face | Real | 1 | 39312 | MOS | 39312 | Overall |
| FIQA (TPAMI2024) [50] | Face | Real | 1 | 42125 | MOS | 42125 | Overall |
| FaceQ | Face | AI-generated, customized and restored | 3 | 12255 | MOS | 32742 | Quality, Authenticity, ID Fidelity, Correspondence |
@@
(1) For face generation models, earlier works such as GANs [17, 31, 81] have demonstrated significant influences. Recently, diffusion models [68, 71, 63] have achieved rapid developments in text-to-image generation. These models can be categorized into pixel-space diffusion models [62, 64, 70, 11], and latent space diffusion models [68, 57, 45, 59, 6]. (2) For face customization, this is a recently defined task that aims to create images of a specific person with the input of an identity reference image and prompts [95]. It can be divided into two categories including conventional test-time optimization [69, 15, 37] and tuning-free customization methods utilizing pre-trained diffusion models [97, 103, 44, 10, 22, 88]. (3) For face restoration, which aims to restore high-quality face images from their low-quality counterparts. Some traditional restoration methods utilize geometric priors [9, 33, 5] or reference priors [41, 42] while in recent years, exploring generative prior such as GAN priors [91, 102, 4, 20, 107] and diffusion priors [94, 84, 49, 104] has become a prevalent trend. Despite these improvements, the generated, customized, and restored images may suffer from various quality issues, highlighting the necessity of evaluators.
@@
Existing face image quality assessment (FIQA) can be divided into Biometric FIQA [1] and Generic FIQA [89, 56, 3, 7]. BFIQA [1, 2, 23, 47, 56] methods are most developed to evaluate the biometric utility of facial components for robust face recognization, which may fail to get satisfying results in terms of perceptual quality [7]. Subsequently, GFIQA methods are designed to improve the face restoration performance and only prioritize the perceptual degradation in face images. Existing FIQA databases are listed in Tab. 1, which predominantly focus on real face images, while the problem of considering face quality in emerging AI-generated images remains unexplored.
@@
In recent years, plenty of researchers have raised interest in assessing AI-generated images(AIGIs) quality [83, 35, 96, 38] by building various databases as shown in Tab 1. These databases can be categorized based on the type of annotation, which includes wide-range pair annotations [35, 96] and MOS annotations [83, 38, 46]. Single-dimensional pair-wise annotations [35] simplify the AIGI quality into an,wu2023humanpreferencescorebetter overall score, which may lead to inaccurate characterization compared to multi-dimensional annatations. All of these databases primarily focus on general AIGI images, leaving a notable gap in datasets dedicated to AI-generated face images.
@@ 3.1.1 FaceQ-Gen Subset Collection
To generate face images, in FaceQ-Gen subset, we design 288 face-centric prompts categorized into nine types: General, Accessories, Clothes, Action, View, Expression, Background, Style, and Facial Attributes. The distribution can be seen in Fig. 1. In particular, for Action categories, we sample from MS-COCO [48] dataset. For other categories, 80 prompts are sourced from [44], while the remaining prompts are self-designed under the assistance of GPT-4o [18]. From the above scheme, we obtain 360 initial prompts as input. A filtering scheme is adopted to guarantee the prompts are face-centric, diverse, and ethical, resulting in 288 final prompts.
Models Collection and Face Image Generation.
The FaceQ-Gen subset comprises 14 recent and representative open-source face generation models, including Deep Floyd [11], Stable Diffusion V1.5 [67], Stable Diffusion V2.1 [67], SD-XL [59], Realistic Vision [65], Stable Cascade [58], ProtoVision V6.6 [60], SD3 [72], Playground V2.5 [39], Kolors [36], PixArt-alpha [6], Hunyuan [45], and Flux-dev [14]. Since our database focuses on the quality assessment of realistic human face images, each model’s negative prompt is configured to exclude “anime” and “semi-realistic” outputs by using terms like “worst quality”, “low quality”, “illustration”, “3D”, “2D”, “painting”, “cartoons”, “sketch”, “anime”, “animation”, “cartoon”, and “semi-realistic”. Safe sensors are enabled to filter out NSFW content. Ultimately, FaceQ-Gen includes a total of 4,032 generated facial images. To ensure diversity in quality, we reduce the sampling steps to one-fourth of each model’s default setting for 111 randomly selected prompts, with all generated images clearly labeled. Please refer to the supplementary material for additional configuration details.
@@ 3.1.2 FaceQ-Cus Subset Collection
Face customization relies on a reference identity image and a guiding prompt. In FaceQ-Cus, the reference images consist of 35 unique identities. Among them, twenty-five identities are sourced from [44], while ten identities are collected by ourselves. The source prompts are selected and filtered from the FaceQ-Gen prompt list, as described in Sec. 3.1.1, with the exclusion of the General class, which may limit customization efficacy. In total, 20 prompts are chosen, and these are reformatted to meet the expected input requirements for different methods by substituting class words or trigger words as needed.
Models Collection and Face Image Customization.
We select six representative identity-oriented image-to-image models for the FaceQ-Cus subset, including ELITE [95], FastComposer [97], IP-Adapter-FaceID [29], InstantID [88], IP-Adapter-Face-ID-Plus [28], and PhotoMaker V2 [44]. Specifically, ELITE [95] requires an additional subject mask, which is obtained using a pre-trained segmentation model [61]. PhotoMaker [44] supports both single and multiple reference images. For a fair comparison, we utilize only a single reference image. Universal negative prompt, NSWF filtering, and face-centric filtering are applied as described in Sec. 3.1.1. Finally, 700 images are generated per method, resulting in a total of 4,200 valid face images. Detailed configurations are provided in supplementary material.
@@ 3.1.3 FaceQ-Res Subset Collection
FaceQ-Res is divided into two categories based on the degradation type of low-quality images: synthetic and real-world. (1) Synthetic: To address the potential underrepresentation in previous works [51, 43], we additionally collect, crop, and align 100 high-quality Asian face images, each with a resolution of 512x512. In addition, we extract key frames from VFHQ [87] to avoid using the training set of restoration models. The data collection pipeline contains blur detection, occlusion detection, keyframe extraction, and keypoint alignment. Samples from the CelebRef dataset are also incorporated. To generate low-quality counterparts, we construct two degraded pipelines following previous works [4, 92, 94, 90] to simulate real-world degradations. In total, we obtain 224 HQ-LQ pairs. (2) Real-world: Real-world low-quality images consist of 223 low-quality images collected from various public datasets, including AgeAB [55], CelebChild [20], Wider [100], LFW [27] and MegaFace [32], ensuring the degradation diversity. Please refer to the supplementary material for more details.
@@
We evaluate several state-of-the-art methods, including diffusion-prior-based methods such as DiffBIR [49], DifFace [104], StableSR [84], as well as GAN- and Transformer-based methods GFPGAN [91], GPEN [102], CodeFormer [107], VQFR [20] and Face-SPARNet [4]. We directly adopt their official code and pre-trained models. Each method is benchmarked on both the F-Res synthetic dataset and the real-world dataset. In total, 4023 restored images are generated, with 2007 images in the real-world case and 2016 images in the synthetic case.
@@ 3.2 Human Preference Dimension Design
Traditional face quality assessments typically rely on a single evaluation dimension, which oversimplifies the evaluation process and fails to capture the unique strengths and weaknesses of different models, as well as individual user preferences. To address these limitations, we introduce a multi-dimensional approach that decomposes “face image quality” into four core dimensions: Quality, Authenticity, ID Fidelity, and Correspondence. These dimensions are further distributed across three categories—face generation, face customization, and face restoration—based on the specific characteristics of each task.
@@
(G-1, C-1, R-1): This dimension evaluates the overall perceptual quality of the image, considering factors such as color accuracy, blur, noise, and artifacts. It is relevant to all three tasks (generation, customization, and restoration) and represents the visual fidelity of the image without considering text alignment or additional reference images.
@@
(G-2): This dimension focuses on how closely the generated image resembles a natural, real-life photograph, with particular emphasis on realistic skin texture, facial details, and wrinkles. It is specifically used in face generation tasks, where ensuring a lifelike appearance is crucial.
@@
(C-2, R-2): This dimension assesses how well the identity of the reference image is preserved in the generated or restored image. It is especially important in face customization and restoration tasks, where maintaining the identity of the individual is paramount.
@@
(G-3, C-3): This dimension evaluates the degree of semantic consistency between the generated or customized image and the provided textual prompt. It is relevant for both face generation and customization tasks, ensuring that the generated image aligns with the specified description.
@@ 3.3 Subjective Experiment
Paticipants and Apparatus.
To ensure a comprehensive, fair, and reliable subjective assessment, 180 participants (90 female and 90 male) are recruited for the experiment, with 60 participants assigned to each task. All participants have normal or corrected-to-normal eyesight. Based on expert feedback, respective tutorials are designed for each task, and all participants complete the tutorial relevant to their assigned task. After the tutorial, participants are instructed to rate 20–30 practice images (10 per dimension) to ensure accurate assessments before proceeding to the main task. For the experimental apparatus, all images are displayed on 27-inch 4K Dell monitors, randomly presented, at their original resolution under standard lighting conditions. The experiment follows the ITU-R BT.500-14 guidelines [73] for subjective evaluations.
@@
We randomly divide 4032 generated images, 4200 edited images, and 4023 restored images into 12 groups, with each group containing approximately 1000 images from a specific task. The subjects are required to rate each image in a group by dragging a sliding window from 0 to 5 with a two-digit decimal on the following 2 or 3 dimensions. (1) Face Generation: Quality and Authenticity are single-stimulus dimensions; the text prompt only appears for the Correspondence dimension. (2) Face Customization: Quality and Correspondence are the same as face generation, while ID Fidelity is double-stimulus, with the ID reference image displayed only when adjusting the ID Fidelity slider. (3) Face Restoration: Quality is single-stimulus, while Fidelity is double- for real-world cases or multi-stimulus for synthetic cases.
Subjective Score Post-processing.
For each group of ratings, outlier detection based on Kurtosis is conducted respectively with a rejection rate at 3% [73]. After removing outliers, the valid raw ratings are converted and scaled into Z-scores ranging from with the following formulas:
@@
where denotes the number of valid subjects, and denotes rescaled z-scores.
Finally, a total of 491,130 reliable raw ratings are collected, with at least 15 valid ratings for each image. In total, 32,742 Mean Opinion Scores (MOS) are calculated across multiple dimensions.

### v2

@@ Abstract
Recent artificial intelligence (AI) generative models have demonstrated remarkable capabilities in image production, and have been widely applied to face image generation, customization, and restoration. However, many AI-generated faces (AIGFs) still suffer from issues such as unique distortions, unrealistic details, and unexpected identity shifts, underscoring the need for a comprehensive quality evaluation method for AIGFs. To this end, we introduce FaceQ, the first comprehensive AI-generated Face image database with fine-grained Quality annotations aligned with human preferences, which consists of 12K images and 491K ratings across multiple dimensions. Using the FaceQ database, we establish F-Bench, a benchmark for comparing and evaluating face generation, customization, and restoration models, highlighting strengths and weaknesses across various prompts and evaluation dimensions. Additionally, we assess the performance of existing image quality assessment (IQA) methods on FaceQ, and further propose a large multimodal model (LMM) based Face quality Evaluator (F-Eval) to accurately assess the multi-dimensional quality of generated faces in a one-for-all manner. Extensive experimental results demonstrate the state-of-the-art performance of our F-Eval. The project page is: https://mediax-sjtu.github.io/F-Bench/.
@@ 1 Introduction
| Database | Domain | Source | Images | Scores | Dimensions |
| AGIQA-3K (TCSVT2024) [51] | General | AI-generated | 2,982 | MOS | Perception, Alignment |
| AIGCIQA (CVPR2024) [89] | General | AI-generated | 2,400 | MOS | Quality, Authenticity, Correspondence |
| RichHF-18K (CVPR2024) [59] | General | AI-generated | 17,760 | MOS | Plausibility, Aesthetics, Text-image Alignment, Overall |
| PIQ23 (CVPR2023) [13] | Face | Real | 5,116 | Pair | Overall, Details, Exposure |
| FaceQ | Face | AI-generated, customized and restored | 12,255 | MOS | Quality, Authenticity, ID Fidelity, Correspondence |
Generative models, including VAE [47], GAN [26], and diffusion models [33, 82, 74], have made remarkable progress in recent years, and have been extended to the fields of text-to-image generation [3], text-driven image editing [44], image restoration [102], etc, leading to high-quality artificial intelligence generated images (AIGIs). As an important aspect of AIGIs, many studies have also investigated face generation [103], face customization [57, 111], and face restoration [62, 90], producing creative, high-quality face images. Despite these advances, many AI-generated faces (AIGFs) still suffer from perceptual deficiencies.
Many natural face image quality assessment databases have been established in the literature [48, 51, 89, 59]. However, natural images are mainly degraded by distortions such as noise, blur, compression [113, 66, 23], etc., while AIGIs generally suffer from unique distortions including unrealistic structures and text-image inconsistencies [59, 89], etc. Recently, many studies have investigated human preferences for AIGIs [59, 89, 50, 39, 106]. However, these general AIGI quality assessment works may not be well aligned with human preferences on AIGFs, since the perceptual quality of AIGFs can be significantly affected by small deviations in facial attributes,(e.g., eye size, skin texture, identity shifts), while the number of AIGFs in these AIGI databases is limited. Moreover, current AIGI quality assessment works mainly focus on the image generation, while human preferences for face customization have rarely been considered, where identity fidelity is significant.
The above limitation highlights the necessity of establishing a comprehensive quality evaluator for AIGFs. To this end, we propose FaceQ, the first AI-generated Face image Quality assessment database with three popular tasks. Specifically, FaceQ contains 12K+ images obtained from 29 generative models across three tasks: 1) face generation, 2) face customization, and 3) face restoration, as shown in Fig.1. Based on the generated images, we collect 491K+ quality ratings from multiple dimensions including quality, authenticity, identity (ID) fidelity, and text-image correspondence, respectively, finally forming 32K+ mean opinion scores (MOSs). Using FaceQ, we establish a human preference benchmark, i.e., F-Bench, for face generation, customization, and restoration models, and thoroughly analyze the generation performance across different evaluation perspectives, models, and categories. In addition, we also compare the correlation of existing image quality assessment (IQA), face quality assessment (FQA), AIGI quality assessment, and preference evaluation metrics with human preferences.
Furthermore, we present F-Eval, the first generated Face image quality Evaluation method designed to perform comprehensive multi-dimensional quality evaluation for generated, customized, and restorted faces in a one-for-all manner. Specifically, F-Eval leverages large multimodal models (LMMs) [11] and adopts instruction tuning [63] for preference-related vision-language adaption. We also apply low-rank adaptation (LoRA) [35] techniques for more efficient feature refining. Extensive experimental results demonstrate that F-Eval outperforms existing quality assessment methods in terms of multiple dimensions relevant to human preference for all three tasks. The main contributions of our work are highlighted as follows:
We construct FaceQ, the first generated face quality assessment database consisting 491K+ human preference ratings from quality, authenticity, ID fidelity, correspondence dimensions, on 12K+ generative face images across three tasks.
We establish F-Bench, a comprehensive human preference benchmark for 29 face generation, customization, and restoration methods. It also extensively evaluates the alignment of 26 existing quality assessment methods with human preferences.
We present F-Eval, a LMM-based quality assessment model to predict multi-dimensional human preference on generated faces. Extensive experiments demonstrate the state-of-art QA performance of the proposed F-Eval.
@@ 2 Related Work
(1) For face generation models, earlier works such as GANs [26, 43, 87] have demonstrated significant influences. Recently, diffusion models [74, 77, 73] have achieved rapid developments in text-to-image generation. These models can be categorized into pixel-space diffusion models [72, 73, 76, 1], and latent space diffusion models [74, 58, 71, 17]. (2) For face customization, this is a recently defined task that aims to create images of a specific person with the input of an identity reference image and prompts [101]. It can be divided into two categories including conventional test-time optimization [75, 24, 49] and tuning-free customization methods utilizing pre-trained diffusion models [104, 111, 57, 20, 30, 94].
(3) For face restoration, which aims to restore high-quality face images from their low-quality counterparts. Some traditional restoration methods utilize geometric priors [19, 46, 16] or reference priors [54, 55] while in recent years, exploring generative prior such as GAN priors [97, 110, 15, 28, 116] and diffusion priors [100, 90, 62, 112, 86, 36, 37] has become a prevalent trend. Despite these improvements, the generated, customized, and restored images may suffer from various quality issues, highlighting the necessity of evaluators.
@@
Existing face image quality assessment (FIQA) can be divided into Biometric FIQA [12, 31, 60, 69] and Generic FIQA [95, 84, 13, 18].
BFIQA methods are most developed to evaluate the biometric utility for face recognization. GFIQA methods [84] focus on perceptual degradation in face restoration task. Both of these databases consist of natural face images, which has distinctly different nature with generated faces.
@@
In recent years, plenty of researchers have raised interest in assessing AI-generated images(AIGIs) quality assessment [89, 48, 51, 23, 107, 40] by building various databases as shown in Tab 1. All of these existing database contain various categories of general AIGI images, leaving a notable gap in datasets dedicated to AI-generated face images.
@@ 3 FaceQ Database Construction
In this section, we introduce the proposed generative face quality assessment database FaceQ. The database includes 12,255 images and 491,130 human opinion ratings. Our FaceQ demonstrates the merits of comprehensive generated visual content and multi-dimensional labeling, which can significantly facilitate the advancement of AIGC research.
@@ 3.1.1 FaceQ-Gen Subset Collection
Models Collection.
We collect 14 representative face generation models, including Deep Floyd [1], Stable Diffusion V1.5 [74], Stable Diffusion V2.1 [74], SD-XL [71], Realistic Vision [8], Stable Cascade [70], ProtoVision V6.6 [7], SD3 [9], Playground V2.5 [52], Kolors [6], PixArt-alpha [17], Hunyuan [58], and Flux-dev [3]. Negative prompts and safe sensors are enabled. In total, FaceQ-Gen includes 4,032 generated face images.
@@ 3.1.2 FaceQ-Cus Subset Collection
Face customization relies on a reference identity image and a guiding prompt. The reference images in FaceQ-Cus are partially self-collected, with the rest sourced from [57]. The prompts are selected and filtered from the FaceQ-Gen prompt list.
Models Collection.
@@ 3.1.3 FaceQ-Res Subset Collection
(1) Synthetic: We construct two degraded pipelines following previous works [15, 98, 100, 96] to simulate real-world degradations on both in-domain data from VFHQ [93], CelebRef [56] and out-of-domain self-collected high-quality Asian face images (2) Real-world: Real-world low-quality images are sourced from various public datasets, including AgeAB [68], CelebChild [28], Wider [108], LFW [38] and MegaFace [45], ensuring the degradation diversity.
@@
We evaluate both diffusion-prior-based methods such as DiffBIR [62], DifFace [112], StableSR [90], and GAN or Transformer-based methods GFPGAN [97], GPEN [110], CodeFormer [116], VQFR [28] and Face-SPARNet [15]. Each method is benchmarked on both the F-Res synthetic dataset and the real-world dataset. In total,FaceQ-Res include 4023 restored images, with 2007 images in the real-world case and 2016 images in the synthetic case.
@@ 3.2 Human Preference Dimension Design
The evaluation dimensions of generated faces need specific design to adequately reflect human preference, given the complexity of perceptual factors involved. To the end, we decompose “quality” into four core dimensions: Quality, Authenticity, ID Fidelity, and Correspondence.
@@
This dimension evaluates the overall perceptual quality of the image, considering factors such as color accuracy, blur, noise, and artifacts. It is used in all tasks.
@@
This dimension evaluates how closely the generated image resembles a natural, real-life photograph, with particular emphasis on realistic skin texture, facial details, and wrinkles. It is used in face generation task.
@@
This dimension evaluates how well the identity of the reference image is preserved in the customized or restored image in face customization and restoration tasks.
@@
This dimension evaluates text-to-image alignment in face generation and customization tasks.
@@ 3.3 Subjective Experiment
To ensure a comprehensive, fair, and reliable subjective assessment, 180 participants with normal or corrected-to-normal eyesight are recruited for the experiment. They first take expert-led tutorials before the main assessment. In the main assessment, participants are asked to rate the image by dragging a sliding window from 0 to 5 with a two-digit decimal from different dimensions. All images are displayed on 27-inch 4K Dell monitors, randomly presented, at their original resolution under standard lighting conditions. The experiment follows the ITU-R BT.500-14 guidelines [78] for subjective evaluations.
Subjective Score Processing.
To eliminate the inter-annotator variability, outlier detection based on Kurtosis is conducted with a rejection rate of 3% [78]. As a result, each image is rated by at least 15 subjects. These valid ratings are converted and scaled into Z-scores ranging from with the following formulas:
@@
where denotes the number of valid subjects for image , and denotes rescaled z-scores.